A company produces optical-fiber cable with a mean of 0.3 flaws per 100 feet. What is the probability that there will be exactly 5 flaws in 1400 feet of cable? Round your answer to four decimal places.
Answer #1

This is a poisson distribution

µ = 0.3 per 100 feet = 0.3 * 14 = 4.2 per 1400 feet

P(X = x) = e * µx / x!

P(X = 5) = e-4.2 * 4.25 / 5! = 0.1633